Give a proof of the theorem, "if n is odd, then n squared is odd".
1
Expert's answer
2021-09-21T03:04:48-0400

Let "n" be an odd number. Then "n=2k+1" for some integer "k." It follows that "n^2=(2k+1)^2=4k^2+4k+1=2(2k^2+2k)+1=2s+1," where "s=2k^2+2k" is an integer number. Therefore, "n" squared is odd.